Science Report – April 17th

Continuing our vegetation ecology and biodiversity project, we plotted transects and surveyed quadrats at two new sights today: Copernicus Valley and Hab Ridge. As yesterday, we surveyed the vegetation along three transects and fifteen quadrats per site, recording abiotic characteristics and the vegetation of each site.

Additionally, we continued to collect voucher specimens for new vascular plant species encountered at each site, adding 15 new collections and ~5 new species for the MDRS operational area.

This evening we plan on entering data and processing data stemming from these surveys.

Operations Report – April 17th

Crew 210 Operations Report 17-APR-2019

SOL: 5

Name of person filing report: Shannon Rupert

Non-nominal systems: nothing to report

Notes on non-nominal systems: nothing to report

Generator (hours run last night): 10 (7:30 am to 9:30 pm)

Solar— SOC 79 % when generator was turned on last night

Diesel Reading – 38 %

Propane Reading – 49 % (Note: I believe I misread this earlier)

Ethanol Free Gasoline – 4 gallons.

Water (auxillary tank) – 0 gallons

Water (static tank) – 300 gallons

Auxillary to Static tank transfer– no

Gallons transferred: No

Water in GreenHab – unknown gallons.

Water (loft) – Static to Loft Pump used – no

Water Meter: not read

Toilet tank emptied: no

Deimos rover used: Still non-operational


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR, yes

Hours: Did not record

Beginning charge: 100

Ending charge: 96

Currently charging: yes

Spirit rover used: yes

Hours: did not check

Beginning charge: 100

Ending charge: 90

Currently charging: yes

Opportunity rover used: yes

Hours: not recorded

Beginning charge: 100

Ending charge: 94

Currently charging: yes

Curiosity rover used: No

Hours: 88.8

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) Honda, 300, 350.1, 350.3

Reason for use: Testing and Field work

Oil Added? No

ATV Fuel Used: 3 Gals.

# Hours the ATVs were Used today:4

Notes on ATVs: Battery in the Honda appears to be going downhill. We
will keep an eye on it and take jumper cables when we use it.

HabCar used and why, where? No

CrewCar used and why, where? Yes, to Hanksville for curry and a test
for live online event tomorrow

General notes and comments: We need a new hose for the compressor.
Current one has several leaks that have been taped.

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: David will send a full report today.

Summary of ScienceDome operations: Things are rockin’ in the lab.
Plant work and soil work in progress.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: First aid kits have been used and
so not all have the necessary supplies to be useful. They should be
gone through and restocked. I will try to get this done before the
next crew arrives.

Questions, concerns and requests to Mission Support: Nothing tonight,
thank you.

Greenhab Report – April 17th

GreenHab Officer: David Murray

Environmental control: Ambient with door open 10:10 to 19:15

Shade Cloth: 80%

Low Temperature: 17 C

High Temperature: 33 C

Relative humidity: 17%

Hours of Supplemental Light: 0

Daily Water Usage for Crops: 20 L total

Water in Blue Tank: very low, needs refilling

Time(s) of Watering: 10:15 and 15:50

Changes to Crops: None.

Narrative: Temperatures were high again today but the GreenHab stayed fairly cool. The fan was running on the highest setting from 11:00 until around 19:00. Crops are looking healthy but I think it’s been too hot for the flowering tomato plant. The fertilizer and lower temperatures from yesterday seemed to help but it’s still showing signs of stress. Harvested parsley and sage that was used in the crew’s lunch and dinner.

Harvest: 21 g parsley, 1 g sage

Sol Summary – April 17th

Crew 210 Sol Summary Report 17-APR-2019

Sol: 5

Summary Title: Plants to the north and plants to the west

Author’s name: Paul Sokoloff

Mission Status: Nominal

Sol Activity Summary: Today we continued our outreach efforts on social media, including answering questions on our Reddit "Ask a Martian Anything", and uploaded a Facebook update video. Four of the crew went to Copernicus Valley to continue transect and quadrat sampling, where we discovered new plant records for the MDRS area, including a stinkweed (Cleomella palmeriana). After a beautiful ride out and back, we returned home to the hab for lunch. In the afternoon we hiked up to the top of Hab Ridge just to the west of MDRS to finish our quadrat sampling. We returned to the hab around 6, and we plan on spending the evening processing samples and pressing plants in the Science Dome.

Look Ahead Plan: Tomorrow AM we will conduct livestream outreach broadcasts (one in French and one in English) in partnership with Discover The Universe and Live It. In the afternoon we will continue to process samples in the science dome, and will meet to discuss next steps for our multi-year program.

Anomalies in work: No problems.

Weather: Bright and Sunny

Crew Physical Status: Good. One crew member slipped into a sinkhole on Sagan Street but caught themselves before falling far, and is OK (one very slight scrape). No problems afterwards

EVA: Four crew members took ATVs to Copernicus Valley in the northwest corner of the MDRS operation area in the morning for quadrat sampling. In the afternoon four crew members hiked up to Hab Ridge for additional quadrat surveys.

Reports to be file: GreenHab Report, Operations Report, Photo Report

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site